Two schools A and B want to award their selected teachers on the values of
honesty, hard work and regularity. The school A wants to award Rs. x each, Rs. y
each and Rs. z each for the three respective values to 3, 2 and 1 teachers with a
total award money of Rs. 1.28 lakhs. School B wants to spend Rs. 1.54 lakhs to
award its 4, 1 and 3 teachers on the respective values (by giving the same award
money for the three values as before). If the total amount of award for one prize on
each value is Rs. 57000, using matrices, find the award money for each value

Grade:upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
The 3 equations are
3x+2y+z=1.28
4x+y+3z=1.54
x+y+z=0.57
\begin{pmatrix} 3& 2 &1 \\ 4& 1 &3 \\ 1 &1 &1 \end{pmatrix}\begin{pmatrix} x\\ y\\ z \end{pmatrix}=\begin{pmatrix} 1.28\\ 0.54\\ 0.57 \end{pmatrix}
|A|=-5 and X=A-1B
A-1=1/5 \begin{pmatrix} 2 &1 &-5 \\ 1 & -2 &5 \\ -3 &1 & 5 \end{pmatrix}
\begin{pmatrix} x\\ y\\ z \end{pmatrix}=1/5 \begin{pmatrix} 2 &1 &-5 \\ 1 & -2 &5 \\ -3 &1 & 5 \end{pmatrix}\begin{pmatrix} 1.28\\ 0.54\\ 0.57 \end{pmatrix}=\begin{pmatrix} 0.25\\ 0.21\\ 0.11 \end{pmatrix}
x=25000,y=21000,z=11000
Thanks and Regards
Shaik Aasif
askIITians faculty